A loan of R100 000, granted at 8% p.a. compounded monthly is to be repaid by regular equal monthly payments over a period of five years, starting one month after the granting of the loan.
The interest portion of the 23rd payment is equal to:
1
Expert's answer
2020-06-02T20:17:36-0400

one payment:

i=0.08/12=0.0066

n=5*12=60

x=(P*i)/((1-(1+i)-n)=(100000*0.0066)/(1-(1+0.0066)-60)=2023.81

principal portion is 1,570.76

 interest portion is 2023.81-1,570.76=453.05